OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x-2y=17 x+3y=17

jimthompson5910 (jim_thompson5910):

Basically they want you to solve for x in the first equation. Then you replace 'x' in the second equation with whatever you got after you solved for x.

OpenStudy (anonymous):

I still don't know how to do that /:

jimthompson5910 (jim_thompson5910):

can you solve x-2y=17 for x?

OpenStudy (anonymous):

x=2y +17 ??

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

now go to the second equation x+3y = 17 and replace 'x' with '2y+17' (since the two are equal) to get (2y+17)+3y = 17 2y+17+3y = 17

OpenStudy (anonymous):

The options are .. (2y+2)+3y=17 (-2y+2)+3y=17 x+3(2y+2)=17 x+3(-2y+2)=17

jimthompson5910 (jim_thompson5910):

So it's (2y+17)+3y = 17, which is shown above

jimthompson5910 (jim_thompson5910):

oh wait, hold on

jimthompson5910 (jim_thompson5910):

hmm there must be a typo in the first equation

jimthompson5910 (jim_thompson5910):

can you double check?

OpenStudy (anonymous):

ooops , My bad . It's x-2y=2

jimthompson5910 (jim_thompson5910):

so x = 2y+2

jimthompson5910 (jim_thompson5910):

and x+3y = 17 becomes (2y+2)+3y = 17

OpenStudy (anonymous):

Thanks & If I was trying to choose a new equation equivalent to y... 15x -y=-6 5x-3y=17 How would I do that ?

jimthompson5910 (jim_thompson5910):

notice we have 15x-y = -6, what do we get when we solve for y?

OpenStudy (anonymous):

y=15x+6 ?

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

Now jump to the other equation 5x-3y=17 and replace y with 15x+6 to get 5x-3(15x+6)=17

OpenStudy (anonymous):

ThankYOU !(:

OpenStudy (anonymous):

yes thats right

jimthompson5910 (jim_thompson5910):

you're welcome